OpenStudy (anonymous):

The length of a rectangle is 4 times its width and the perimeter is 150 feet. What is the width of the rectangle?

OpenStudy (anonymous):

This one is just a little complicated. Please help!

OpenStudy (photon336):

what does 4 times the length mean to you?

OpenStudy (photon336):

@PhoenixSoul

OpenStudy (anonymous):

Solve the following for w, the width: 2 (w + 4 w) = 150

OpenStudy (trojanpoem):

Let the width be x length = 4 width = 4x  Perimeter = sum of all sides x + 4x + 4x + x = 150 10x = 150 x = 15

OpenStudy (photon336):

should really go step by step

OpenStudy (nuttyliaczar):

Set up an equation to represent this scenario. L=4W, and P=150=2(L+W)=2(4W+W)=2(5W)

OpenStudy (anonymous):

ok thanks a bunch :) I will write this down in my notes for later.

OpenStudy (anonymous):

yes it is 15 for the width srry